# LightMDETR: A Lightweight Approach for Low-Cost Open-Vocabulary Object Detection Training

### v1

@@ IV-B Dowstream Tasks
LightMDETR-CF 86.80 88.76 81.78 79.10 84.12 71.07 81.06 80.81
TABLE II: Accuracy performance comparison between our proposed models, LightMDETR and LightMDETR-CF, and other detection models in the referring expression comprehension task on the RefCOCO, RefCOCO+, and RefCOCOg datasets. For testing, RefCOCO and RefCOCO+ datasets are evaluated using person vs. object splits: ”testA” includes images with multiple people, while ”testB” includes images with multiple objects from other categories. RefCOCOg features two distinct data partitions.
Table II presents a comparison of our models, LightMDETR and LightMDETR-CF, against other detection models on RefCOCO, RefCOCO+, and RefCOCOg. RefCOCO and RefCOCO+ are evaluated using person vs. object splits: ”testA” includes images with multiple people, while ”testB” includes those with multiple objects. There is no overlap between training, validation, and testing images. RefCOCOg is split into two partitions.
@@
LightMDETR-CF 85.37 95.52 96.73 77.98 93.85 95.47 80.24 94.26 96.56
TABLE III: Precision performance comparison between our proposed models, LightMDETR and LightMDETR-CF, and MDETR in the referring expression comprehension task on the RefCOCO, RefCOCO+, and RefCOCOg datasets.
Results presented in Table III showcase the precision performance of our models, LightMDETR and LightMDETR-CF, in comparison to MDETR on the RefCOCO, RefCOCO+, and RefCOCOg datasets. Precision at rank $k$ (P@k) indicates the percentage of correct predictions within the top $k$ ranked results. Specifically, P@1 measures precision at the top-1 prediction, P@5 within the top 5, and P@10 within the top 10.
Our models demonstrate competitive performance, with LightMDETR achieving the highest precision at P@1 on RefCOCO (85.92%) and RefCOCOg (80.97%), surpassing MDETR slightly on these datasets. Furthermore, LightMDETR-CF leads in P@5 on RefCOCO (95.52%) and P@10 on RefCOCOg (96.56%), highlighting the effectiveness of our lightweight approach. Although MDETR performs marginally better on RefCOCO+, LightMDETR closely follows, validating our hypothesis that freezing the backbone and training only the DFE component allows our models to achieve comparable, if not superior, performance with reduced computational complexity. This confirms the efficiency and accuracy of our proposed models in the referring expression comprehension task.
@@ V Conclusion
We propose a novel method for training open-vocabulary object detection models that significantly reduces the number of parameters to tune. Our approach leverages specialized pretrained encoders for text and images, which remain frozen during training. The only component we train is a lightweight module we developed called the ”Deep Fusion Encoder” (DFE). The DFE is designed to encode features from both the text and image encoders using shared parameters. To enable this, we introduce a learnable parameter called ”context,” which identifies the source of each feature. This context is embedded in the DFE representation, allowing it to seamlessly switch between processing text and image features.
Our method, when integrated into the MDETR model, outperforms the baseline in terms of accuracy and precision on the referring expression comprehension tasks across the RefCOCO, RefCOCO+, and RefCOCOg datasets. Importantly, this approach is not limited to MDETR; it can be applied to any open-vocabulary object detection model to reduce training costs while maintaining high performance.
For future work, we plan to conduct further experiments to validate our method on tasks such as phrase grounding, referring expression segmentation, and visual question answering. Additionally, we will explore its application to other open-vocabulary object detection models.

### v3

@@ IV-B2 Referring expression comprehension
LightMDETR-Plus 86.80 88.76 81.78 79.10 84.12 71.07 81.06 80.81
TABLE II: Accuracy performance comparison between our proposed models, LightMDETR and LightMDETR-Plus, and other detection models in the referring expression comprehension task on the RefCOCO, RefCOCO+, and RefCOCOg datasets. For testing, RefCOCO and RefCOCO+ datasets are evaluated using person vs. object splits: ”testA” includes images with multiple people, while ”testB” includes images with multiple objects from other categories. RefCOCOg features two distinct data partitions.
Table II presents a comparison of our models, LightMDETR and LightMDETR-Plus, against other detection models on RefCOCO, RefCOCO+, and RefCOCOg. RefCOCO and RefCOCO+ are evaluated using person vs. object splits: ”testA” includes images with multiple people, while ”testB” includes those with multiple objects. There is no overlap between training, validation, and testing images. RefCOCOg is split into two partitions.
@@
LightMDETR-Plus 85.37 95.52 96.73 77.98 93.85 95.47 80.24 94.26 96.56
TABLE III: Precision performance comparison between our proposed models, LightMDETR and LightMDETR-Plus, and MDETR in the referring expression comprehension task on the RefCOCO, RefCOCO+, and RefCOCOg datasets.
Results presented in Table III showcase the precision performance of our models, LightMDETR and LightMDETR-Plus, in comparison to MDETR on the RefCOCO, RefCOCO+, and RefCOCOg datasets. Precision at rank $k$ (P@k) indicates the percentage of correct predictions within the top $k$ ranked results. Specifically, P@1 measures precision at the top-1 prediction, P@5 within the top 5, and P@10 within the top 10.
Our models demonstrate competitive performance, with LightMDETR achieving the highest precision at P@1 on RefCOCO (85.92%) and RefCOCOg (80.97%), surpassing MDETR slightly on these datasets. Furthermore, LightMDETR-Plus leads in P@5 on RefCOCO (95.52%) and P@10 on RefCOCOg (96.56%), highlighting the effectiveness of our lightweight approach. Although MDETR performs marginally better on RefCOCO+, LightMDETR closely follows, validating our hypothesis that freezing the backbone and training only the UP module allows our models to achieve comparable, if not superior, performance with reduced computational complexity.
Figure 6: An illustration of LightMDETR on segmentation with the model fine-tuned on phraseCut.
IV-B3 Referring expression segmentation
Referring expression segmentation involves pinpointing and delineating objects in images using textual cues, as demonstrated with the PhraseCut dataset [37]. This dataset features images sourced from VG, complete with segmentation masks for a variety of expressions, many of which refer to multiple objects. Following the approach of MDETR, our training unfolds in two phases. Initially, we fine-tune our pre-trained model for 10 epochs while keeping ResNet-101 and RoBERTa frozen, optimizing for precise bounding box predictions and employing box AP for early stopping. In the subsequent phase, we freeze the network weights and focus on training a segmentation head for 35 epochs, implementing a learning rate adjustment at 25 epochs, supervised by a blend of Dice/F1 loss [38] and Focal loss [39]. During inference, we assign each predicted box a confidence score of $1-P(\emptyset)$, filtering out those below a threshold of 0.7. Ultimately, we consolidate the masks from the selected boxes into a unified binary mask corresponding to the referring expression.
Method M-IoU Pr@0.5 Pr@0.7 Pr@0.9
RMI [40] 21.1 22.0 11.6 1.5
HULANet [31] 41.3 42.4 27.0 5.7
MDETR 53.1 56.1 38.9 11.9
LightMDETR 53.45 56.98 39.12 11.6
LightMDETR-Plus 53.87 57.07 39.27 11.82
TABLE IV: Validation of Referring Expression Segmentation using the mean intersection-over-union (IoU) between predicted and ground-truth masks, alongside precision Pr@$I$, where success is defined as the predicted mask achieving an IoU with the ground-truth that exceeds the threshold $I$.
The results in Table IV highlight the effectiveness of our proposed methods, LightMDETR and its enhanced variant LightMDETR-Plus. Both methods demonstrate superior performance compared to MDETR, achieving a mean intersection-over-union (M-IoU) of 53.45 and 53.87, respectively. Notably, they also exhibit improved precision at various thresholds, particularly at Pr@0.5 and Pr@0.7, with LightMDETR-Plus leading the metrics.
Downstream tasks such as phrase grounding, referring expression comprehension, and segmentation demonstrate that our proposed lightweight framework significantly enhances the efficiency of open-vocabulary object detection training. By considerably reducing the number of trainable parameters, it maintains or even improves performance on these tasks as illustrated in Fig. 5 and 6.
@@ V Conclusion
In this paper, we present a novel method for training open-vocabulary object detection systems that significantly reduces the number of parameters to tune. Our approach utilizes specialized pre-trained encoders for text and images, which remain frozen during training. The only trainable component is a lightweight module we developed, termed the ”Universal Projection” (UP) module. This module efficiently encodes features from both text and image encoders using shared parameters. To facilitate this, we introduce a learnable parameter known as the ”modality token”, which identifies the source of each feature. This modality token is integrated into the UP representation, allowing for seamless transitions between processing text and image features.
When implemented within the MDETR model, our method demonstrates superior accuracy and precision across tasks such as phrase grounding, referring expression comprehension, and segmentation. Importantly, this approach is not restricted to MDETR; it can be adapted as a modular framework for any open-vocabulary object detection system, effectively reducing training costs while preserving high performance.
Looking ahead, we aim to explore the application of this lightweight framework to other open-vocabulary object detection systems, further enhancing their efficiency and effectiveness.
Acknowledgments
We would like to express our deepest gratitude to LabCom IRISER for their financial support, which made this work possible as part of an internship. Their commitment to fostering research and innovation has been invaluable throughout this project. We are extremely thankful for the opportunity and the resources they provided, enabling us to pursue and complete this work successfully.